Brush-Script-Italic Fonts

Maurizio Loreti

Rev. 1.0 — Feb. 11, 1995

The files here included allow to use a PostScript Type-1 font named “Brush-Script-Italic” in your TEX documents: the important ones are pbsi.pfa (the font in encoded ASCII form) and pbsi0.afm (the font metric file), while the others have been generated with the fontinst package by Alan Jeffrey and are included here only for sake of simplicity. What you need to do in order to install this font is:

1. to copy the .tfm (font metric) and the .vf (virtual font) files in the appropriate TEX font directories;

2. to copy the .fd (font definition) files in the TEX input directory;

3. to copy the .pfa file where your DVI to PS converter will look for it; if you are using dvips, you want to modify the file psfonts.map adding a

     line saying “pbsi0 <pbsi.pfa”.

4. All of my TEXing is done on VAX/VMS and Amigas, but I have access to a complete TEX setup on VAXen only; on the Amiga, TEX and dvips process my files correctly, but the previever refuses to work :-(. I hope you will have no problems with different environments.

Some short notes: I have found the original Brush-Script-Italic font on the Internet, as a NeXT screen font; and I have heavily modified the original file on my Amiga personal computer, with a font drawing package named TypeSmith. My modified version (as well as the original font) are both public domain; you are free to modify my files as you want, as long as you don’t make any profit from them.

You will find here a file called sample.tex that is an example of the actual use of this new font; it is a real document I have typeset (in Italian, sorry; but you can see a couple of accented characters correctly generated by the T1 font encoding). Be careful, do not use TEX commands that change the typeface (like \bfseries or \sffamily); Brush-Script-Italic has no defined variants!

Happy TEXing! Maurizio Loreti

                                      EMail: loreti@padova.infn.it
                   WWW: http://mvxpd5.pd.infn.it/wwwcdf/mlo.html